### v2

@@ Introduction
If the network has some degree of asymmetry, we could proceed identically as for undirected networks. The only difference would be the usage of the expressions for directed networks in order to obtain the elements $p_{k}$ of the process $\mathcal{P}$ (see Supplementary Material for the full derivation)
@@
However, contraty to the undirected case, although $p_{k}$ reduces to the same expression in Eq. 10 if the modularity matrix is symmetric, the first element of the process does not coincide with the modularity function to optimize for the full graph (i.e., $p_{0}\neq Q(C^{0})$). This is undesirable (but not critical) because it would mean that the first step of the hierarchy needs to be separated from the rest. Furthermore, the symbolic expression for the corresponding QUBO problem would possess $N(N+1)$ symbolic terms, potentially compromising the computational workflow for large graphs. Lastly, the $ij$-th and $ji$-th entries would be associated with the product of binary variables $x_{i}x_{j}$. Then, the QPU would interpret the coupling of the two qubits as a sum of $B_{ij}^{d,\text{k}}+B_{ji}^{d,\text{k}}$, hence loosing all the non-symmetric information.
To address this, the two off-diagonal entries should be combined into a single descriptive interaction instead of two separate contributions. To leverage the desirable properties of the process $\mathcal{P}$, we adopt an elegant solution from Leicht and Newman [28]. Notably, $p_{k}$ in Eq. 9 is a scalar, so its transpose $p_{k}^{\top}$ has the same global minima. Therefore, for directed networks, the linear combination $\frac{1}{2}\left(p_{k}+p_{k}^{\top}\right)$ can be optimized recursively instead of optimizing $p_{k}$ alone. In practice, this modification updates the generalized modularity matrix as follows:
@@
Importantly, all the desirable properties showed for the undirected scenario, hold true for the equation above: 1) if $k=0$, the element $\frac{1}{2}\left(p_{k}+p_{k}^{\top}\right)$ reduces to the QUBO in Eq. 5; 2) $\text{if }x_{i}=x_{j}\ \ \forall i,j\in k\subset V\implies p_{k}=0$; and 3) for undirected networks, this symmetrized version of the generalized modularity matrix reduces to Eq. 11. Recall that points 1 and 2 ensure that we remain within a QUBO formalism without linear terms and that the first binary division is obtained by optimizing the correct function. The latter point presents a convenient form from the software perspective.
Algorithmic details
@@
As shown in pseudocode for the hierarchical annealing process in Algorithm 1, we have used a common recursive approach to divide and conquer types of algorithms. First, for each network, we computed a generalized modularity matrix (lines 3-4) and built the appropriate QUBO (line 5). Next, we sampled results from the quantum annealing process for that QUBO (line 6). Unless stated otherwise in the manuscript, to indicate optimal binary community division we chose the best result from $100$ different samples (i.e. numreads parameter of the Advantage_system5.4 version of a D-Wave sampler). Next, we recursively called the same procedure for both returned subcommunities unless one of them is empty (lines 7-9). Otherwise, we returned the non-empty part of the division (lines 10-13).
The optimal solution, in terms of the modularity index, the community structure, and the total computational time used, was defined as the solution that yielded the highest modularity index (see Algorithm 2). For that, we ran the hierarchical process 20 times independently. In each one of these runs, the returned modularity was kept only if its value was higher than the previously assigned one (line 7). The D-Wave Advantage sampler was initialized on every one of these runs. The elapsed time was measured from the start of the hierarchical process until the last subdivision was attained; not including the computation of the modularity matrices, the QUBO functions, or the embedding times. It should be noted that the time we report is highly dependent on the specific configuration of the local machine. Therefore, exact quantitative analyses of this variable may not be directly interpretable; nonetheless, the relative times for different runs possess valuable information. We provide a general description of the software package together with snippets of the high-level Python code developed in the Supplementary Material, enhancing the usability of quantum annealing resources for optimization purposes.
Algorithm 1 Recursive QA for modularity maximization
1:Initialize graph $G=(V,E)$
2:procedure Binary__QA($G$, $V$)
3:     Compute modularity matrix B (Eq. 4, undirected) or Bd (Eq. 6, directed)
4:     Compute the generalized modularity matrix Bk in Eq. 13
5:     QUBO $\leftarrow$ Compute the $k$-th element of $\mathcal{P}$ in Eq. 10
6:     $k_{1}$, $k_{2}\leftarrow$ QA binary optimization of the QUBO
7:     if $k_{1}\neq\varnothing$ and $k_{2}\neq\varnothing$ then
8:         Binary__QA($G$, $k_{1}$)
9:         Binary__QA($G$, $k_{2}$)
10:     else if $k_{1}\neq\varnothing$ then
11:         return $k_{1}$
12:     else
13:         return $k_{2}$
@@
15:end procedure
Algorithm 2 Modularity maximization
1:Initialize graph $G=(V,E)$
2:m $=0$, $C=\varnothing$, dendrogram =absent=$\varnothing$= ∅
[…]
4:     $p_{k}$, $k$, $D\leftarrow$ Binary__QA($G$, $V$)
5:     Compute Q in Eq. 1 from $k$
[…]
7:         m $\leftarrow$ Q
8:         $C\leftarrow k$
9:         dendrogram $\leftarrow D$
10:     end if
11:end for
Technical aspects of the D-Wave minor embedding
The QUBO problem, illustrated by Eq. 10, can be depicted as a weighted (and/or directed) graph. In this representation, the vertices correspond to binary variables $x_{i}$, and the edges signify the quadratic terms $x_{i}x_{j}$ in the problem formulation. Linear coefficients are attributed as weights to the vertices, while quadratic coefficients are similarly assigned to the edges.
@@
The actual D-Wave QPU topology uses a Pegasus graph of degree 15 [29]. The QUBO graph must be mapped onto that architecture, which is achieved by minor-embedding [30] i.e. mapping each logical binary variable to a group of physical qubits on the actual machine so all required connections are realized. As the QUBO function based on the generalized modularity matrix (see Eq. 10) is dense, we used complete graph embedding predefined routine (clique embedding). To reduce the time of the embedding procedure, we used the cache mechanism to compute embedding for all necessary clique sizes only once and reuse it while performing the hierarchical approach described by Algorithm 1.
Other community detection algorithms and their multiple solutions
To compare the performance of the hierarchical process and pure quantum annealing, we evaluated results using four different approaches: the Louvain, Leiden, Bayan, and Hierarchical Gurobi algorithms, as well as a hybrid quantum-classical method based on discrete quadratic optimization models (DQM) [10, 12, 14, 25]. For the Leiden and Louvain algorithms, we computed multiple instances for each network. Specifically, each algorithm was run a certain number of times per network, and the community structure with the highest modularity was retained as the “true” structure. The DQM method demonstrated high stability, likely due to its problem partitioning strategy for fitting into the annealer [24, 25]. The Hierarchical Gurobi used the same algorithms as the Hierarchical annealing but only used the classical Gurobi solver. For the Bayan, which relies heavily on the Gurobi optimizer and its corresponding license, the Louvain, and Leiden algorithms we ran 20 instances per network. The Bayan algorithm is known to return exact solutions for networks up to a considerable size [15].
@@
Community detection algorithms often yield varying structures due to the use of heuristics or inherent solution degeneracy [31]. We evaluated the variability between communities identified by two algorithms through 1) consensus classification matrices and 2) the Dice-Sørensen score. For each network, an $N\times N$ consensus coclassification matrix recorded how often nodes $i$ and $j$ were grouped into the same community across algorithm runs [32]. Perfect agreement between algorithms would result in matrix entries $M_{ij}$ of 0 or 1 (or 0% and 100%), indicating stable or well-defined community structures. Conversely, a range of intermediate $M_{ij}$ values suggests inconsistent community assignments. While informative, this matrix does not assess community structure optimality, necessitating the additional use of a modularity index. To compare community structures directly, we computed the Dice score between pairs of communities as:
@@
We measured the improvement of the classical algorithm w.r.t. the Hierarchical annealing by computing the relative increase. For that, we extracted the maximum modularity from the pool of solutions (see 2), and computed the relative error percentage-wise. Thus,
%$\%RI=100\cdot\frac{\text{Max }Q_{HA}-\text{Max }Q_{Z}}{Q_{Z}},$ (15)
@@
Network generation, power-law exponent $\alpha$, and scale-free networks
We generated random networks using three different models. First, the Erdős-Rényi model [33] connected any pair of initially disconnected nodes with a constant probability $p$, producing random networks without clear statistical organization. Despite this, non-trivial geometrical behaviors can still emerge, making it a useful null model for study. Second, the Barabasi-Albert model [34] added new nodes dynamically, where the probability of a new node $i$ connecting to node $j$ was proportional to $j$’s degree, $p_{j}\propto k_{j}$. This preferential attachment mechanism produced networks with robust scaling properties and enhanced resilience. Finally, we used the Holme-Kim model [35], which incorporates preferential attachment but adds a triad formation step to boost clustering, defined as the presence of connected triangles in the network. This produced highly clustered, power-law networks. To test our approach in directed networks, we also used the preferential attachment mechanism applied to directed versions of the graphs [36]. This yielded power-law and scale-free networks with directed edges. The parameters of this model were left as the default values, both in Bollobás, et al., 2003 [36], and in the NetworkX Python package [37].
For all of these networks, we fitted a discrete power-law distribution to their node degrees present in a given network [38]. A network is said to have a power-law distribution or structure if the probability $P(k)$ of a given node to have degree $k$ follows an exponential distribution,
$P(k)\sim k^{-\alpha},$
where $\alpha$ is the power-law exponent. Furthermore, if this exponent lies within 2 and 3, the network exhibits a scale-free distribution, where the probability of highly connected nodes does not vanish even for high degrees $k$ [34]. The study of complex networks with scale-free properties is of crucial because they are ubiquitous in social, technological, and biological environments.
Results
Hierarchical annealing in simple networks
As a first benchmark, we generated chains of 3-cliques for undirected and directed examples and treated each one of these 3-cliques as proxies for true communities. Hierarchical annealing and Louvain algorithms were executed 5 times per network, and their performance was compared based on returned modularities (Fig. 1a), number of found communities (Fig. 1b), and variability of the algorithms in terms of their consensus classification matrices (Fig. 1c-d; see Methods). For these experiments, the resolution parameter $\gamma$ was left to 1. For networks with up to 30 nodes, both algorithms achieved identical modularities in all 5 runs. The hierarchical annealing algorithm maintained stable and optimal performances as the network size increased, achieving the same maximum modularity as Louvain in all but one network. Notably, it even surpassed Louvain’s maximum modularity by 0.5% for a network of 150 nodes (Fig. 1a). Interestingly, for a network of 60 nodes, the annealing algorithm produced a slightly less optimal solution while demonstrating no variability. This suggests that the inability to explore the state space of possible solutions may lead to suboptimal behaviors. Overall, the annealing algorithm exhibited slightly less variability than Louvain without compromising solution optimality.
We show a graphic example in Fig. 1f for an undirected and directed network of 8 3-cliques (i.e., 24 nodes) where, despite finding slightly different partitions, the modularities are identical (i.e., 0.625 and 0.6295 for the undirected and directed chains respectively).

### v3

@@ 2.4 Directed networks
If the network has some degree of asymmetry, we could proceed identically as for undirected networks. The only difference would be the usage of the expressions for directed networks to obtain the elements $p_{k}$ of the process $\mathcal{P}$ (Supplementary Material S2)
@@
However, contrary to the undirected case, the first element of the process does not coincide with the modularity function to optimize for the full graph (i.e., $p_{1}\neq Q(C^{2})$). This is undesirable because it would mean that the first step of the hierarchy needs to be separated from the rest. Furthermore, the $ij$-th and $ji$-th entries would be associated with the product of binary variables $x_{i}x_{j}$. Then, the QPU would interpret the coupling of the two qubits as a sum of $B_{ij}^{d,\text{k}}+B_{ji}^{d,\text{k}}$, hence losing all the non-symmetric information.
To address this, the two off-diagonal entries should be combined into a single descriptive interaction instead of two separate contributions. To leverage the desirable properties of the process $\mathcal{P}$, we adopt an elegant solution from Leicht and Newman [Leicht2008]. Notably, $p_{k}$ in Eq. 9 is a scalar, so its transpose $p_{k}^{\top}$ has the same global minima. Therefore, for directed networks, the linear combination $\frac{1}{2}\left(p_{k}+p_{k}^{\top}\right)$ can be optimized recursively instead of optimizing $p_{k}$ alone. In practice, this modification updates the generalized modularity matrix as follows:
@@
Importantly, all the desirable properties showed for the undirected scenario hold: 1) if $k=1$, the element $\frac{1}{2}\left(p_{1}+p_{1}^{\top}\right)$ reduces to the QUBO in Eq. 5; 2) $\text{if }x_{i}=x_{j}\ \ \forall i,j\in k\implies p_{k}=0$; and 3) for undirected networks, this symmetrized version of the generalized modularity matrix reduces to Eq. 11. Points 1 and 2 ensure that we remain within a QUBO formalism without linear terms and that the first binary division is obtained by optimizing the correct function. The latter point presents a convenient form from the software perspective.
2.5 Algorithmic details
We ran the hierarchical process $N_{runs}=20$ times independently (see Algorithm 1). Since, in practice, reaching $\max Q$ once is enough, this number ensures a competitive performance across all networks tested in this work (Fig. S2; see also Supplementary Material S3). More specifically, modeling each of the 20 outcomes as a single Bernoulli event, we observed probabilities of success higher than 40%, and, in most cases, $\max Q$ was found with frequencies narrowing 100%. In general, however, users should tailor this limit based on available computational resources and the behavior of the specific problem instance. As a rule of thumb, more runs translate into a more extensive exploration of the annealer’s variability. In each of these runs, the returned modularity was kept only if its value was higher than the previously assigned one (line 11). The D-Wave Advantage sampler was initialized on every one of these runs. We provide a general description of the software package together with snippets of the high-level Python code developed in the Supplementary Material, enhancing the usability of quantum annealing resources for optimization purposes [lamza_qhyper_2024].
@@ 2.5 Algorithmic details
As shown in the pseudocode for the Hierarchical annealing process in Algorithm 1, we have used a common recursive approach to divide and conquer types of algorithms. First, for each network, we computed a generalized modularity matrix (lines 20-21) and built the appropriate QUBO (line 22) and loaded the corresponding clique embedding from the cache (line 23). Next, we sampled results from the quantum annealing process for that QUBO (line 24-26). In the Advantage_system5.4 version of a D-Wave sampler, we chose the best result from $100$ different samples (i.e., num_reads=100). The annealing_time was set to $20.00\pm 0.01\ \mu s$, and the rest of the parameters were also left to their respective default values. Next, we recursively called the same procedure for both returned subcommunities unless one of them was empty (lines 27-34).
Algorithm 1 Hierarchical quantum annealing (QA) for modularity maximization
[…]
2:Compute and cache clique embedding of size $\forall n\leq|V|$
[…]
5:m =0=0, $C=\varnothing$, dendrogram $\varnothing$
6:
7:# Repeated modularity maximization
8:B or Bd $\leftarrow$ Modularity matrix (Eqs. 4, 6)
[…]
10:  $C^{k}$, $D$ $\leftarrow$ Binary__QA(B or Bd, $V$, $\gamma$)
11:  Compute $Q$ in Eq. 1 from $C^{k}$
[…]
13:   m, $C$, dendrogam $\leftarrow$ $Q$, $C^{k}$, $D$
@@
15:end for
16:
17:# Recursive quantum annealing routine
18:procedure Binary__QA(B, $V$, $\gamma$)
19:  BVV $\leftarrow$ Generalized Mod. matrix (Eq. 13)
20:  QUBO $\leftarrow$ $k$-th element of $\mathcal{P}$ (Eq. 10)
21:  QUBOE $\leftarrow$ Load clique embedding of size $|V|$
22:  $k_{1}$, $k_{2}\leftarrow$ QA(QUBOE, num_reads, annealing_time)
23:  if $k_{1}\neq\varnothing$ and $k_{2}\neq\varnothing$ then
24:   Binary__QA(B, $k_{1}$, $\gamma$)
25:   Binary__QA(B, $k_{2}$, $\gamma$)
26:  else if $^{k}_{1}\neq\varnothing$ then
27:   return $k_{1}$
28:  else
29:   return $k_{2}$
30:  end if
31:end procedure
2.6 Technical aspects of the D-Wave minor embedding
@@
The Advantage_system5.4 D-Wave QPU topology uses a Pegasus graph of degree 16. The QUBO graph must be mapped onto that architecture, which is achieved by minor-embedding, i.e., mapping each logical binary variable to a group of physical qubits on the actual machine so all required connections are realized. As the QUBOs are dense, we utilized a predefined routine for complete graph embeddings (clique embedding). To reduce the time, we cached the embeddings for all clique sizes only once and reused them accordingly in Algorithm 1. Crucially, although Hierarchical annealing relies on pure quantum processes, the embedding remains a classical step. The chain strength was set using D-Wave’s default strategy, which determines the value based on the number of logical qubits and the variance in connectivity of the physical qubits. Broken chains were resolved using majority voting, assigning each logical qubit the most common spin value along its chain.
2.7 Computational complexity and running times
For a problem of size $n$, the Hierarchical annealing procedure comprises two main stages: (i) computation of the modularity matrix and (ii) recursive binary QA optimization. The first stage is executed only once before entering the recursion. Using big-$\mathcal{O}$ notation, scales as $\mathcal{O}(n^{2})$ due to an outer product between degree vectors. The second stage requires a more nuanced analysis, as each recursion step invokes a QA optimization.
At each step of the recursion, the generalized modularity matrix needs to be computed, followed by a QA instance. The former is done in linear time $\mathcal{O}(n)$, while the latter depends on the specific problem to be solved. For convenience, we denote the computational time required to perform a single binary QA step as $f(n)$, and the total accumulated annealing time as $\mathcal{K}_{n}$. Recall that the final complexity should consider B, thus yielding $\mathcal{O}(n^{2}+\mathcal{T}(n))$ where $\mathcal{T}(n)$ is the complexity of the recursion tree.
Certain instances of the recursion tree grow linearly with the size of the network $n$, but these are uncommon as they imply that each subdivision yields isolated nodes. Instead, a more reasonable outcome is a dynamic and unbalanced recursive tree for which the Akra-Bazzi theorem provides an asymptotic solution to the expected computational complexity (Supplementary Material S4). Briefly, the theorem states that recurrences of the form
$\mathcal{T}(n)=\sum_{i=1}^{s}a_{i}\mathcal{T}(b_{i}n+h_{i}(n))+g(n),$ (14)
have an asymptotic solution,
$\mathcal{T}(n)\sim n^{p}\left(1+\int_{n_{0}}^{n}\frac{g(u)}{u^{p+1}}du\right).$ (15)
The parameter $p$ satisfies the characteristic equation $\sum_{i=1}^{s}a_{i}b_{i}^{p}-1=0,$ with $a_{i}>0$, $0<b_{i}<1$ for all $i=1,\ldots,s$, and $g(n)$ is a continuous, positive, and regular function for all sufficiently large $n$. The functions $h_{i}(n)$ represent small perturbations in the arguments.
In general, a community $k$ is split into two communities $k_{1},k_{2}$ of arbitrary sizes $|k_{1}|=n_{k}/a_{k}$ and $|k_{2}|=n_{k}(1-1/a_{k})$, where $n_{k}$ is the number of nodes in community $k$ and $1/a_{k}$ is the fraction of nodes in $k$ assigned to $k_{1}$. The ratio $a_{k}$ will differ for every community $k$, but it can be modeled as a random number: $a_{k}=a+\delta_{k}+O(\delta_{k}^{2})$ with $\delta_{k}\sim U(-a,L)$, and $1\leq a\leq L$ for some finite $L$ depending on the average ratio along the tree. Then, the community sizes can be expressed as,
$\begin{split}&\frac{n_{k}}{a_{k}}=\frac{n_{k}}{a}\left(\frac{1}{1+\delta_{k}/a}\right)=\frac{n_{k}}{a}+h_{1}(n_{k})\\ &n_{k}\left(1-\frac{1}{a_{k}}\right)=n_{k}\left(1-\frac{1}{a}\frac{1}{(1+\delta_{k}/a)}\right)=n_{k}\left(1-\frac{1}{a}\right)+h_{2}(n_{k})\end{split}$ (16)
where $h_{1}(n_{k})$ and $h_{2}(n_{k})$ represent bounded perturbations stemming from variable community sizes and rounding errors (e.g., floor and ceiling functions $\cdot$]). Additionally, the resolution limit guarantees that the recursion will halt no later than at some minimum effective size $n_{0}\in[1,n-a]$, where $\mathcal{T}(n_{0})=\mathcal{O}(1)$.
Formally, the time to compute the whole recursion for a community $k$ is a function of the two attached sub-problems and the corresponding operations taking place at the same level,
$\small\begin{split}\mathcal{T}(n_{k})=\qquad&\mathcal{T}\left(\left[\frac{n_{k}}{a}\right]+h_{1}(n_{k})\right)+\\ &\qquad\mathcal{T}\left(\left[n_{k}\left(1-\frac{1}{a}\right)\right]+h_{2}(n_{k})\right)+n+f(n),\end{split}$ (17)
which satisfies Eq. 14 if $s=2$, $a_{1}=a_{2}=1$, $b_{1}=\frac{1}{a}$, $b_{2}=1-\frac{1}{a}$, $g(n)=n+f(n)$, and $p=1$ (see Supplementary Material). Crucially, the QA complexity $f(n)$ seems exponential at worst [Albash2018], which still satisfies the conditions of regularity. Solving for Eq. 15 gives a complexity of $\mathcal{T}(n)\sim\mathcal{O}(n\log n+\mathcal{K}_{n})$ with
$\mathcal{K}_{n}\doteq n\int_{n_{0}}^{n}\frac{f(u)}{u^{2}}du.$ (18)
While $f$ is generally unknown, it can be empirically measured and compared with theoretical assumptions (Table 1). Lastly, while the resolution parameter $\gamma$ influences tree depth through the resolution limit, it does not change the computational complexity of the recursion. However, it might impact the annealing times because it modifies the underlying Hamiltonian [Reichardt2006, Traag2011].
$f(n)$ $\mathcal{K}_{n}$ Total complexity $\mathcal{T}(n)$
$c$ $O(1)$ $\mathcal{O}(n\log n)$
$\log n$ $O(\log n)$ $\mathcal{O}(n\log n+\log n)\sim\mathcal{O}(n\log n)$
$n$ $O(n\log n)$ $\mathcal{O}(n\log n)$
$n^{s},\ s>1$ $O(n^{s})$ $\mathcal{O}(n\log n+n^{s})\sim\mathcal{O}(n^{s})$
$e^{n^{s}},\ s\in\mathbb{R}^{+}$ n.c.f. $\mathcal{O}(ne^{n^{s}})$
Table 1: Asymptotic behavior of $\mathcal{K}_{n}$ and $\mathcal{T}(n)$ for different $f(n)$ growth rates. The parameter $c$ is any positive and finite constant. ’n.c.f.’ stands for No Closed Form (see section S4.3 in the Supplementary Material).
The total execution time comprised of four components: (i) local operations, such as creation of Python objects; (ii) retrieval of cached embeddings; (iii) network communication with the Advantage solver; and (iv) service time on the QPU [^1] https://docs.dwavequantum.com/en/latest/quantum_research/operation_timing.html. The first is implementation-dependent and not intrinsic to the algorithm. The second can be measured using canonical Python tools and, albeit dependent on hardware specifics, could depend non-trivially on the problem’s size and topology. Network communication includes problem transmission and result retrieval, whereas the service time comprises queueing, preprocessing, and QPU execution. For user convenience, network communication and service are provided jointly as a single Sample function in the DWAVE API, a name we use throughout the rest of the paper for readability. Although the QPU access time accounts for all numreads, this definition reflects practical usage, as obtaining a viable solution typically requires multiple samplings.
We evaluated the scalability of the Hierarchical annealing procedure by computing the maximum modularity for networks of increasing size and varying resolution parameters, recording the corresponding elapsed times. For each experiment, we collected three metrics: QPU access, cache-read embedding, and combined sampling–communication time (which included QPU access but excluded embedding retrieval). Measurements were obtained at each recursion step and aggregated to compute the total execution time per problem. Because Algorithm 1 was run iteratively, we considered the running times of the solution that yielded the maximum modularity.
Lastly, because paid usage agreements typically govern access to quantum computing resources, we also quantified the official computing time consumed in the D-Wave Leap Dashboard. This enabled a fair comparison between standard D-Wave approaches, such as discrete quadratic models (DQM), and our custom implementation, providing practical guidance for estimating resource consumption when planning D-Wave computational grants. We manually recorded the available usage time immediately before and after each full execution of the modularity maximization routine, as well as for the DQM runs (Fig. S3).
2.8 Network generation, power-law exponent, scale-free networks, and stochastic block models
We generated random networks using three different models. First, the Erdős-Rényi model [erdds1959random] connected any pair of initially disconnected nodes with a constant probability $p$, producing random networks without clear statistical organization. Despite this, non-trivial geometrical behaviors can still emerge, making it a useful null model for study. Second, the Barabasi-Albert model [Barabasi1999] added new nodes dynamically, where the probability of a new node $i$ connecting to node $j$ was proportional to $j$’s degree, $p_{j}\propto k_{j}$. This preferential attachment mechanism produced networks with robust scaling properties and enhanced resilience. The Holme-Kim model [Holme2002] incorporated preferential attachment and a triad formation step to boost clustering. This produced highly clustered, power-law networks. To test our approach in directed networks, we also used the preferential attachment mechanism applied to directed versions of the graphs [bollobas2003]. This yielded power-law and scale-free networks with directed edges. The parameters of this model were left as the default values, both in Bollobás, et al., 2003 [bollobas2003], and in the NetworkX Python package [hagberg2008].
We fitted a discrete power-law distribution to their node degrees present in a given network [alstott2014powerlaw]. A network is said to have a power-law distribution or structure if the probability $P(k)$ of a given node having degree $k$ follows an exponential distribution,
$P(k)\sim k^{-\alpha},$
where $\alpha$ is the power-law exponent. Furthermore, if this exponent lies between 2 and 3, the network exhibits a scale-free distribution, where the probability of highly connected nodes does not vanish even for high degrees $k$ [Barabasi1999].
Lastly, we used the stochastic block model (SBM) to evaluate the Hierarchical annealing procedure against known ground truths [abbe2018]. Networks consisted of $N=100$ nodes assigned to three planted communities of sizes 50, 40, and 10. Following [Karrer2011], nodes belonging to communities $r,s$ were connected with a probability
$\omega_{rs}=\lambda\omega^{\text{planted}}_{rs}+(1-\lambda)\xi,$ (19)
where $\lambda\in[0,1]$ is a mixing parameter, $\xi\sim U(0,1)$, and
$\boldsymbol{\omega}^{\text{planted}}=\begin{pmatrix}p_{in}&p_{out}&p_{out}\\ p_{out}&p_{in}&p_{out}\\ p_{out}&p_{out}&p_{in}\end{pmatrix}$. (20)
$p_{in}$ and $p_{out}=\tfrac{1}{2}(1-p_{in})$ controlled the intra- and inter-community connection probabilities, respectively. We generated 10 networks for each $\{p_{in},\lambda\}$ pair, with $p_{in}\in\{0.6,0.7,0.8,0.9\}$ and $\lambda\in[0,0.9]$ ins steps of 0.1. For $\lambda=0$, this is equivalent to the planted partition model, enabling a principled comparison between modularity maximization and probabilistic inference methods [Newman2016], with task difficulty increasing monotonically with $\lambda$.
2.9 Other algorithms and multiple solutions
We evaluated results using several classical algorithms: Louvain [Blondel2008], Leiden [Traag2019], Bayan [aref2022] — known to return exact solutions for networks up to a considerable size [aref2023], Hierarchical Gurobi, spectral clustering [Damle2018], simulated annealing [see code], Infomap [Rosvall2008], and degree-corrected stochastic block model inference [Karrer2011, Peixoto2014]; as well as DQM [Wierzbinski2023], a hybrid quantum-classical method. Each algorithm was run a certain number of times per network, and the community structure with the highest modularity was retained as the ”true” structure.
Unless stated otherwise, Louvain, Leiden, Infomap, spectral clustering, and simulated annealing were each run 100 times per network, whereas SBM inference was run 1000 times to ensure. Hierarchical Gurobi followed the same recursion as Hierarchical annealing but used the classical Gurobi solver; both it and Bayan were run for 20 instances due to licensing constraints. DQM was run five times, as its solutions were highly stable. Both DQM and simulated annealing required a predefined maximum number of communities: for DQM, we used the resolution limit ($\max|C|\to\sqrt{n}+1$), while for simulated annealing, we used the number of communities found by Hierarchical annealing, as performance otherwise degraded to near random. For simulated annealing, we used T_max=10, T_min=0.001, and 50 updates.
@@ 2.9 Other algorithms and multiple solutions
Community detection algorithms often yield varying structures due to the use of heuristics or inherent solution degeneracy [fornito2016fundamentals]. We evaluated the variability between communities identified by two algorithms through 1) consensus classification matrices and 2) the Dice-Sørensen score. For each network, an $N\times N$ consensus coclassification matrix recorded how often nodes $i$ and $j$ were grouped into the same community across algorithm runs [Lancichinetti2012]. Perfect agreement between algorithms would result in matrix entries $M_{ij}$ of 0 or 1 (or 0% and 100%), indicating stable or well-defined community structures. Conversely, a range of intermediate $M_{ij}$ values suggests inconsistent community assignments. While informative, this matrix does not assess community structure optimality, necessitating the additional use of a modularity index. To compare community structures directly, we computed the Dice score between pairs of communities as:
@@
We measured the improvement of the classical algorithm w.r.t. the Hierarchical annealing by computing the relative increase. For that, we extracted the maximum modularity from the pool of solutions and computed the relative error percentage-wise. Thus,
$\%RI=100\cdot\frac{\text{Max }Q_{HA}-\text{Max }Q_{Z}}{Q_{Z}},$ (22)
@@
When node assignments were pre-defined, we compared the recovered communities to the true structure using the variation of information, split–join distance, normalized mutual information, and the adjusted Rand index; all implemented in igraph [csardi2006igraph].
3 Results
3.1 Hierarchical annealing in simple networks
As a first benchmark, we generated chains of 3-cliques for both undirected and directed examples, treating each one of these 3-cliques as a proxy for true communities. Hierarchical annealing and Louvain algorithms were executed 5 times per network, and their performance was compared based on returned modularities (Fig. 1a), number of found communities (Fig. 1b), and variability of the algorithms in terms of their consensus classification matrices (Fig. 1c-d; see Methods). For these experiments, the resolution parameter $\gamma$ was left to 1. For networks with up to 30 nodes, both algorithms achieved identical modularities in all 5 runs. The Hierarchical annealing algorithm maintained stable and optimal performances as the network size increased, achieving the same maximum modularity as Louvain in all but one network. Notably, it even surpassed Louvain’s maximum modularity by 0.5% for a network of 150 nodes (Fig. 1a). Interestingly, for a network of 60 nodes, the annealing algorithm produced a slightly less optimal solution, yet demonstrated no variability. This suggests that the inability to explore the state space of possible solutions may lead to suboptimal behaviors. Overall, the annealing algorithm exhibited slightly less variability than Louvain without compromising solution optimality.
We show a graphic example in Fig. 1f for an undirected and directed network of 8 3-cliques (i.e., 24 nodes) where, despite finding slightly different partitions, the modularities are identical (i.e., 0.625 and 0.6295 for the undirected and directed chains, respectively). Similar results hold for a chain of 1-cliques (i.e., nodes), where the absence of any recursive structure does not impede the algorithm from obtaining high modularity indices (Fig. S4).